OpenStudy (anonymous):

Police can estimate the speed of a vehicle before the brakes are applied using the formula , 0.75d= s^2/30.25 where s is the speed in miles per hour and d is the length of the vehicle’s skid marks. What was the approximate speed of a vehicle that left a skid mark measuring 160 feet?

OpenStudy (anonymous):

.75(160)=S^2/30.25

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yeah, i plugged 160 in for d. can you solve for s?

OpenStudy (anonymous):

i wasn't sure that's why i asked:/

OpenStudy (anonymous):

120=S^2/30.25 3630=S^2 S=?

OpenStudy (anonymous):

1815?

OpenStudy (anonymous):

No... \[s=\sqrt{3630}\]

OpenStudy (anonymous):

60 approximatley

OpenStudy (anonymous):

Yep